# On Efficient Variants of Segment Anything Model: A Survey

### v2

@@ 4.1 Datasets and Metrics
We utilize COCO 2017 [74] and LVIS v1 [45] as the evaluation datasets. COCO is a large-scale dataset for object detection, segmentation, and captioning, containing 330K images and 1.5 million object instances across 80 object categories. LVIS is tailored for large vocabulary instance segmentation, featuring over 2 million high-quality segmentation masks across more than 1200 categories in 164K images. For our evaluation, we use the validation sets of both datasets, which include 36,781 instances in 5,000 images for COCO and 244,707 instances in 19,809 images for LVIS. To evaluate efficiency, we first test several soft indicators like #$\#Params$, $FLOPs$, and $Memory\ Usage$. And we further calculate the Efficient Error Rate ($EER$) for a more comprehensive evaluation, as outlined in [98]. $EER$ is defined as,
@@
where $N$ is the number of metrics and $M_{i}$, $R_{i}$ refer to the i-th metric of the tested model and the reference one, respectively. In our evaluation, we set the reference model to be SAM-H. In addition to these metrics, we also report the runtime and throughput of these models. For accuracy evaluation, we use mean intersection over union ($mIoU$) for the SegAny task and mean average precision ($AP$) for instance segmentation.
4.2 Efficiency Comparison
@@ 4.3 Accuracy Comparison
For instance segmentation tasks, we adopt the ViTDet [66], YOLOv8 [60], GrounddingDINO [79], Detic [177] and H-Deformable-DETR [59] with Swin-L [81] as object detectors to generate bounding boxes of latent objects, with reference to the experiments conducted in [129, 176, 171]. We evaluate the average precision (AP) of all objects, as well as AP of small, medium, and large objects. The results are reported in Tab. 6, 7 and 8. Similar to the previous results, we find that on COCO dataset EfficientViT-SAM-XL1 always presents the highest AP with any of the detectors (except H-Deformable-DETR). Under the setting of equipping ViTDet as detector and testing on LVIS dataset, SAMfast-H surpass all other variants with AP of 44.5%.
@@
Based on the results in Section4.2 and Section 4.3, we further make a throughput-mIoU scatter to observe variants’ efficiency-accuracy trade-off. Specifically, we select the throughput and mIoU that are evaluated on COCO dataset with ground truth bounding boxes as prompts. The results are illustrated in Fig. 16.

### v3

@@ 4.1 Datasets and Metrics
We utilize two fundamental datasets, COCO 2017 [219] and LVIS v1 [220], and two more challenging benchmarks, SGinW [^1] https://eval.ai/web/challenges/challenge-page/1931/overview and UVO v1.0 [221], for comprehensive evaluation. COCO is a large-scale dataset for object detection, segmentation, and captioning, containing 330K images and 1.5 million object instances across 80 object categories. LVIS is tailored for large vocabulary instance segmentation, featuring over 2 million high-quality segmentation masks across more than 1200 categories in 164K images. For our evaluation, we use the validation sets of both datasets, which include 36,781 instances in 5,000 images for COCO and 244,707 instances in 19,809 images for LVIS.
Segmentation in the Wild (SGinW) benchmark consists of 25 different segmentation datasets, spanning a wide range of real-world domains, from common objects like fruits and animals to rare concepts like rail and brain tumor[222]. Unidentified Video Objects (UVO) is a challenging benchmark for dense, open-world segmentation in videos, with numerous ”unseen” objects not belonging to any of the COCO categories or the LVIS categories. It is worth noting that, for UVO benchmark, we only leverage the FrameSet for image-level evaluation, which includes 8124 frames sampled from 2708 videos with 65588 annotated instances.
To evaluate efficiency, we first test several soft indicators like #$\#Params$, $FLOPs$, and $Memory\ Usage$. And we further calculate the Efficient Error Rate ($EER$) for a more comprehensive evaluation, as outlined in [55], to measure how much the variants are more efficient in comparison with the vanilla SAM-H. Specifically, $EER$ is calculated by averaging the ratios of a set of efficient metrics between evaluated models and the reference model. The definition of $EER$ is illustrated as follows,
@@
where $N$ is the number of metrics and $M_{i}$, $R_{i}$ refer to the i-th metric of the tested model and the reference one, respectively. Consequently, tested models with higher efficiency will achieve lower $EER$. In our evaluation, we select {#$\#Params$, $FLOPs$} as metrics and set the reference model to be SAM-H. In addition to these metrics, we also report the runtime and throughput of these models. For accuracy evaluation, we use mean intersection over union ($mIoU$) for the SegAny task and mean average precision ($AP$) for instance segmentation.
@@
[…]
**footnotetext: The results of FastSAM [41] are collected from [46] and others are evaluated under our framework.
4.2 Efficiency Comparison
[…]
@@ 4.3 Accuracy Comparison
For instance segmentation tasks, we adopt the ViTDet [223], YOLOv8 [169], GrounddingDINO [224], Detic [225] and H-Deformable-DETR [226] with Swin-L [227] as object detectors to generate bounding boxes of latent objects, with reference to the experiments conducted in [113, 46, 122]. We evaluate the average precision (AP) of all objects, as well as AP of small, medium, and large objects. The results are reported in Tab. 7, 8 and 9. Similar to the previous results, we find that on COCO dataset EfficientViT-SAM-XL1 always presents the highest AP with any of the detectors. Under the setting of equipping ViTDet as detector and testing on LVIS dataset, SAMfast-H surpass all other variants with AP of 44.5%.
To further evaluate variants’ segmentation ability in more challenging scenarios, we conduct instance segmentation on SGinW and UVO benchmark, equipped with GrounddingDINO for box prompts generation by following [104]. Results are illustrated in Tab. 10 and Tab. 11, respectively. On SGinW benchmark, we observe that the rewritten SAMfast-H achieves the highest mean AP over 25 datasets, with only 0.1% decrease compared to SAM-H, followed by EfficientViT-SAM-XL1 and EfficientSAM-S with 0.4% and 0.8% more drop respectively. From the perspective of dataset, the segmentation results on House-Parts, Brain-Tumor and Rail are significantly lower than others, indicating them more challenging datasets. On UVO benchmark, we notice that EfficientViT-SAM-XL1 demonstrates the best performance with AP of 31%, exceeding SAM-H by 1.1%, while SAM2.1-B+ also surpasses original SAM by 1.0%.
Table 10: Quantitative results of zero-shot instance segmentation on SGinW benchmark with GroundingDINO as the object detector.We report variants’ Average Precision (AP) on each dataset and mean AP over all 25 datasets.
Dataset
SAM-H
SAMfast-H
sam2-B+
FastSAM
MobileSAM
EdgeSAM
EfficientSAM-Ti
EfficientSAM-S
RepViT-SAM
SlimSAM50
SlimSAM77
TinySAM
Q-TinySAM
EfficientViT-SAM-L0
EfficientViT-SAM-XL1
NanoSAM
Airplane-Parts 35.3 35.2 36.6 23.1 31.1 31.6 32.5 35.1 29.6 29.1 26.1 28.6 33.0 35.2 34.6 30.5
Bottles 65.7 65.2 66.5 59.6 64.1 64.9 63.4 64.8 65.1 63.8 64.0 64.6 63.1 66.5 67.1 63.7
Brain-Tumor 11.7 11.7 11.2 2.1 12.2 11.3 12.0 12.0 11.6 11.2 11.6 9.9 11.8 11.5 12.1 11.9
Chicken 85.9 86.8 79.1 71.1 84.5 81.9 83.4 85.6 85.2 83.2 85.5 83.2 82.7 82.4 85.1 80.2
Cows 46.3 46.2 46.6 17.0 42.7 40.1 43.3 45.3 42.2 43.0 41.5 38.2 41.4 43.5 45.3 40.5
Electric-Shaver 70.5 70.2 71.8 49.4 66.3 65.2 66.2 68.5 64.2 58.1 63.9 59.9 65.6 69.8 70.6 64.2
Elephants 77.2 77.3 76.2 61.3 70.5 68.4 70.7 74.1 70.0 68.6 70.1 64.3 68.4 72.2 74.8 68.1
Fruits 82.3 82.3 80.7 81.5 82.3 81.5 81.7 81.7 81.5 81.5 81.5 74.0 76.2 81.5 81.5 82.3
Garbage 23.7 23.9 25.9 10.1 22.7 20.5 19.7 23.3 21.9 17.8 19.2 16.9 21.8 23.5 23.2 20.7
Ginger-Garlic 45.8 45.8 44.9 28.3 43.6 40.2 44.6 45.7 47.0 39.5 40.8 39.4 46.1 43.9 46.7 42.6
Hand 76.6 76.5 78.2 81.6 73.2 71.0 69.3 76.7 73.0 71.9 74.6 66.9 70.8 74.1 75.3 70.7
Hand-Metal 81.4 81.1 78.8 70.6 80.9 78.5 76.4 80.0 81.3 77.8 78.8 78.7 78.4 79.2 79.6 78.5
HouseHold-Items 60.1 60.1 60.1 60.1 58.8 58.8 58.8 60.1 60.1 60.1 60.1 58.8 59.5 60.1 60.1 58.8
House-Parts 8.4 8.3 8.0 6.6 8.1 8.1 8.0 8.3 8.3 8.1 8.0 7.8 7.9 8.2 8.3 8.1
Nutterfly-Squireel 70.0 69.4 80.8 23.1 63.1 57.5 68.5 70.8 59.7 52.8 53.7 59.5 60.4 63.5 67.8 60.3
Phones 34.9 35.0 30.2 28.4 33.9 33.4 34.0 35.5 34.5 33.8 35.3 34.2 34.5 34.3 34.8 31.4
Poles 23.3 23.3 15.6 5.1 16.6 15.9 20.0 27.7 13.4 14.2 12.5 4.7 12.5 19.0 16.7 6.9
Puppies 50.1 50.1 50.1 28.4 45.0 44.6 46.7 47.8 45.0 46.0 46.8 38.6 44.2 45.8 50.1 43.8
Rail 15.8 15.7 12.2 1.0 14.1 13.5 12.9 13.4 12.9 12.4 15.1 17.5 15.4 17.5 16.7 13.1
Salmon-Fillet 36.2 36.3 19.7 24.5 29.9 31.7 12.4 21.8 25.2 27.8 28.0 18.3 27.0 39.2 37.1 25.6
Strawberry 83.5 83.3 83.6 79.8 83.3 83.0 82.7 83.3 83.6 82.6 82.4 83.7 82.6 83.4 83.6 83.5
Tablets 29.6 29.5 27.1 29.0 28.7 28.4 29.0 29.3 28.9 28.7 28.7 28.6 28.5 28.8 28.8 28.7
Toolkits 20.9 20.9 20.2 12.3 19.3 18.6 19.1 20.7 18.5 17.7 18.6 19.2 18.6 19.9 20.6 19.1
Trash 29.7 29.7 27.7 27.3 28.8 28.3 28.8 29.1 28.9 28.6 28.7 27.5 28.2 28.7 29.3 28.6
Watermelon 64.9 64.5 63.7 44.4 63.5 66.2 64.6 66.1 62.7 62.1 63.7 62.8 63.4 67.3 67.0 63.2
Mean AP 49.2 49.1 47.8 37.0 46.7 45.7 45.9 48.3 46.2 44.8 45.6 43.4 45.7 48.0 48.7 45.0
Table 11: Quantitative results of zero-shot instance segmentation on UVO benchmark with GroundingDINO as the object detector.
Model $\rm AP$ $\rm AP^{S}$ $\rm AP^{M}$ $\rm AP^{L}$
SAM-H 29.9 10 20.8 44.9
SAMfast-H 29.7 9.9 20.7 44.6
SAM2-B+ 30.9 9.4 21.3 47
FastSAM 20.8 7 14.7 30.1
MobileSAM 25.2 8.2 17.4 38
EdgeSAM 24.9 8.6 17.9 36.4
EfficientSAM-Ti 25.3 7.5 17.9 38.1
EfficientSAM-S 28 9.5 19.3 41.9
RepVit-SAM 26.2 8.9 18.2 39.1
EfficientViT-SAM-L0 29.5 9.2 19.9 45.1
EfficientViT-SAM-XL1 31 9.8 21.1 47.2
SlimSAM-50 26.2 9.2 18.6 38.6
SlimSAM-77 25.1 8.7 17.8 37
TinySAM 22.9 8.1 16.5 33.4
Q-TinySAM 24.8 8 17.6 37
NanoSAM 22.9 6.2 15.4 35.5
@@
Based on the results in Section 4.2 and Section 4.3, we further make a throughput-mIoU scatter to observe variants’ efficiency-accuracy trade-off. Specifically, we select the throughput and mIoU that are evaluated on COCO dataset with ground truth bounding boxes as prompts. The results are illustrated in Fig. 18. For most of the variants, the improvement in latency accompanies with the degradation in accuracy. For instance, NanoSAM at the bottom-right corner, demonstrates the highest throughput and lowest mIoU, while models around the top-left corner present poorer efficiency and better accuracy. Besides, it is also worth noting that, EfficientViT-SAM-L0 on the top-right exhibits both the highest mIoU and relatively high throughput, achieving best efficiency-accuracy trade-off among these variants.
@@ 5 Conclusion
Data Availability. The data supporting the experiments of this study are openly available, as stated in Section 4.1.